James researched these nutrition facts:
1 g of soy milk has 0.005 g of carbohydrates and 0.030 g of protein.
1 g of vegtables has 0.14 g of carbohydrates and 0.030 g of protein.
James wants his lunch to have 50 g of carbohydrates and 20 g of protein.
how many grams of soy milk and vegtables does he need?

Answer by josgarithmetic(39625)

x for grams of milk(soy) and y for grams of vegetables.

Account for carbohydrates:(0.005)x+(0.14)y=50
5x+140y=50000
x+28y=10000

Account for protein: (0.030)x+(0.030)y=20
3x+3y=2000

You may want to try adjustment of these for use of the Elimination Method, the system being:
----------------------
x+28y=50000
----------------------
3x+3y=2000
----------------------


3x+84y=300000
-
3x+3y=2000

81y=300000-2000=298000
y=298000/81
y=3679 grams of vegetables (probably wrong)

x+28y=50000
x=50000-28y
x=50000-28(3679)
x=-53012 Which obviously makes no sense.

Either I made a mistake which I have not found (after rechecking the whole solution once) or one of your given datum is wrong. This would otherwise be a reasonably simple two-variable linear system exercise.